OpenStudy (kingtazz99):

The area of a rectangle is 99 in². The length of the rectangle is 2 in. longer than the width. What is the width of the rectangle?

OpenStudy (stamp):

area = length *width, or a = lw area is given A =99, length is given l = 2 +w, and width is simply w. substituting we realize: a = lw 99 = (2 +w) w 99 = 2w +ww ww +2w -99 = 0 Solve w

OpenStudy (mathmale):

Your input? People here will guide you towards finding your own answers. Please don't expect to be given answers outright.